Bondi Corporation makes automotive engines. For the most recent month, budgeted production was 2,600 engines. The standard power cost is $1.70 per machine-hour. The company's standards indicate that each engine requires 10.4 machine-hours. Actual production was 2,900 engines. Actual machine-hours were 28,890 machine-hours. Actual power cost totaled $53,702.

Required:

Determine the rate and efficiency variances for the variable overhead item power cost. (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Input all amounts as positive values.)
Answer #1

a

Variable overhead rate variance

$4589

U

b

Variable overhead efficiency variance

$2159

F

Part A

Variable overhead rate variance = (AH*AR)-(AH*SR) = 53702-(28890*1.70) = $4589 Unfavorable

Part B

Variable overhead efficiency variance = SR* (AH-SH) = 1.70*(28890-(2900*10.40)) =1.70*(28890-30160) = -2159 = $2159 Favorable